Write the equation of the line that passes through the point (2,-4) and is perpendicular to the line whose equation is y= -1/2x + 3
Please HELP!!

Answer by scott8148(6628)

perpendicular lines have slopes that are negative reciprocals of each other

so, the slope of the new line is -1/(-1/2) or 2

using point-slope __ y-(-4)=2(x-2) __ y+4=2x-4 __ y=2x-8